Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Tuesday, December 01, 1998

The following is a brief summary of the environmental monitoring activities that were completed on December 01, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

12/01/98 (Tuesday)

Segment 1 report for 12/1/98: Due to heavy rain most crews were unable to work today. The hydrostatic testing crew worked on welding test heads for the segment to be tested from EMP 25.43 to 21.75. Shoring was being installed in the bore pit on the north side of Stream Crossing 9 (EMP 15.35). At Emidio Pump Station, work on the office interior continued.

Segment 2B (Castaic to Sylmar): Report for 12-1-98. Drilling crews at Feedmill Rd and the Santa Clara river completed pulling the reamer and anticipate drawing the 6" sleeve in the morning. At Calgrove Blvd, crews pulled a shoring box at the southern bore pit and completed the backfilling of the trench. South of Weldon Cyn Rd, trenching crews neared EMP 71.7 with pipe and welding crews nearing EMP 71.9. Due to rain crews limited welding and coating activities. Slurry and fiber optic conduit was completed to EMP 72.34 while 200' of asphalt was completed to EMP 72.5. Segment 2C (Templin Hwy to Castaic): Due to recent rain events, reveg. efforts have sprouted within the right of way.

Segment 3A, 3B, & 3C (Sylmar to Los Angeles): Report for 1 December 1998: 3A: Rain fell throughout the day, but did not halt construction activities. Backfill crews focused on closing up the ditch around Tujunga Wash today. Those crews also did minor work at Hollywood Way. The bore was short a few feet from completion at the Tuxford bore site. Welding, x-ray, and coating focused on Pacoima Wash. The pipe was placed on the hangers at Pacoima Wash. 3B: Electrical work continued at the valve at Cornfield Yard, ~MP 98.48. 3C: Welding, x-ray, and coating moved to work on Segment 5. Backfill worked on pouring slurry in the road near MP 73.86, as well as clean up. Work on the cover at the high speed channel continued.

Inspection Summary

In the past 24 hours, 3 inspection event(s) were reported for 3 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page